March 23, 1998

The L^{A}TEX2_{ϵ} Seminar Group Department of Mathematics University of South Carolina Columbia, SC 29208

George F. McNulty Noble Prize Committee Royal Society of Sweden Stockholm, Sweden

Dear George,

This is the body of the letter. It can contain both text (like this) and mathematics. The mathematics can be in-line x^{2} + y^{2} = z^{2} or displayed

\frac{∂u}{∂t} = △u + f(u).

You can even put a theorm-like environment in the letter.

Lemma 1. The letter style will be used by many members of the Department.

This should be enough to demonstrate that (essentially) anything can be put in the body of a letter. Sincerely, George McNulty

Sincerely,

George McNulty Professor